OpenStudy (anonymous):

lim (x,y)-->(0,0) (x+y)^2/(3x^2+2y^2) along the y-axis

OpenStudy (anonymous):

even along the y axis?

OpenStudy (anonymous):

i believe you take the partial derivative of that function with respect to x and y so it will be df/dx=2x/6x=2/6...using l'hospitals rule the answer along x axis is 2/6 df/dx=2y/4y=1/2...limit is 1/2 along the y-axis

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

what if the denominator were 3x^2 +4y^2? this time along the x-axis, would it be 2x/6x ?